The frequency of a sound wave is 4000hz and its wavelength is 17.5cm. How much time wiil it take to cover a distance of 3.5km?


1
Expert's answer
2019-09-16T10:00:06-0400

The speed of a sound wave

"v=f\lambda=4000\:\rm Hz\times 0.175\: m=700\: m/s"


So, the time needed for wave to cover a distance 3.5 km


"t=\frac{d}{v}=\frac{3500\:\rm m}{700\:\rm m/s}=5\: \rm s"